OpenStudy (anonymous):

Find the area of a square whose perimeter is 64m?

OpenStudy (anonymous):

64 divided by the number of sides to find one side first

mathslover (mathslover):

do you know what is the formula for area and perimeter of a square ?

OpenStudy (anonymous):

perimeter is 64 so x+x+x+x=64 4x=64 x=16 now area is area=x^2 area=16^=256 m^2